OpenStudy (cutiecomittee123):

which of these are correct pi/6 is a reference angle to 5pi/6 3pi/6 8pi/6 13pi/6 I know that 5pi/6 is one of them

OpenStudy (anonymous):

\[\frac{ 13 \pi }{ 6 }\]

OpenStudy (anonymous):

\[\frac{ 13 \pi }{ 6 }=2\pi+\frac{ \pi }{ 6 }\]

OpenStudy (cutiecomittee123):

why dont the others work

OpenStudy (cutiecomittee123):

and do any of the others work

OpenStudy (anonymous):



OpenStudy (cutiecomittee123):

okay so then the only two that can fit in are 5pi/6 and 13pi/6

OpenStudy (cutiecomittee123):

are you sure because i know that it can go into other quadrants as a reference angle as well

OpenStudy (anonymous):



OpenStudy (anonymous):

pi-5pi/6=pi/6

OpenStudy (anonymous):

yes you are correct.

OpenStudy (anonymous):

it can go in other quadrants also. so 5pi/6 and 13pi/6